Following particulars has been shared for you. If fixed cost is Rs500000, selling price is Rs60 and Variable cost is Rs20

a. Calculate and describe -Breakeven point (5 Marks)

b. What could be the sales number to earn a profit of Rs 50000, at a Fixed Cost Rs 200000, Variable Cost Rs30 per unit, Selling Price Rs 60 per unit


1
Expert's answer
2021-10-03T14:14:23-0400

a. The breakeven point is:

BEP = 500,000/(60 - 20) = 12,500 units.

b. The sales number to earn a profit of Rs 50000, at a Fixed Cost Rs 200000, Variable Cost Rs30 per unit, Selling Price Rs 60 per unit is:

TR - TC = TP,

60Q - 200,000 - 30Q = 50,000,

30Q = 250,000,

Q = 8,333.33 units.